A = 8kg , 10 ms-1 and moves to East

B = 5kg , 20ms-1 and moves to North

-Before Hitting-


A+B = 13kg
& -After Hitting-


if the hitting is fully unelastic, find the velocity of A and B after hitting.What is the direction that A+B moves after hitting.
1
Expert's answer
2011-07-14T06:34:35-0400
We should add the& vectors& of impulse A and B
Pa = MaVa = 8*10 = 80 kg*m/s (to the East)
Pb = MbVb& = 5*20 = 100 kg m/s (to the North)

The modulus of the impulse after inelastic hitting is
radic; (Pa2 + Pb2) = raidc; (802 + 1002) = 128.1 kg m/s
The speed :
128.1 / (8 + 5) = 9.85 m/s
tan (angle to the North) = 80/100 = 0.8
38.66 degrees from North to East.